How would i solve these two word problems???

1)An express and local train leave GraysLake at 3 P.M. and head for Chicago 50 miles away. The express travels twice as fast as the local and arrives 1 hour ahead of the local. Find the speed of each train.
2)Joe has a collection of nickels and dimes that is worth $5.65. If the number of dimes were doubled and the number of nickels were increased by 8, the value of the coins would be $10.45. How many dimes does he have?


Answer by stanbon(48568)

2)An express and local train leave GraysLake at 3 P.M. and head for Chicago 50 miles away.
The express travels twice as fast as the local and arrives 1 hour ahead of the local. Find the speed of each train.
-----------
Express DATA:
Distance = 50 miles : Rate = 2x mph ; time = 50/2x
---------
Local train DATA:
Distance = 50 miles ; Rate = x mph ; time = 50/x
-----------------
EQUATION:
local train time - express train time = 1 hr
50/x - 50/2x = 1
Multiply thru by 2x to get:
100 - 50 = 2x
50 = 2x
x = 25 mph ( speed of the local train)
2x = 50 mph ( speed of the express)
-----------------------------------------------
Joe has a collection of nickels and dimes that is worth $5.65.
If the number of dimes were doubled and the number of nickels were increased by 8, the value of the coins would be $10.45.
How many dimes does he have?
----------------
EQUATIONS:
5n + 10d = 565 cents
5(n+8) + 10(2d) = 1045
--------------------------
5n + 10d = 565
5n + 20d = 1005
-----------------------
Subtract 1st from 2nd to get:
10d = 440
d = 44 (# of dimes)
------------
Substitute into 5n+10d=565 to solve for "n"
5n+10*44 = 565
5n + 440 = 565
5n = 125
n = 25 (# of nickels)
=============
cheers,
Stan H.